OpenStudy (anonymous):

Need help! During the last 10 years, the population of a town increased from 2,000 to 8,320. What percent increased does this represent? Show work please

OpenStudy (anonymous):

Ii need help with numbers 5 and 16

OpenStudy (tkhunny):

2000/2000 = 1.0000 = 100% This is a 0% increase 4000/2000 = 2.0000 = 200% This is a 100% increase 8000/2000 = 4.0000 = 400% This is a 300% increase

OpenStudy (anonymous):

I'm confused.

OpenStudy (tkhunny):

I'm just trying to point out that there is a little disconnect between a simple ratio and a "Percent Increase". Just create the simple ratio and then think about the question, in light of my little demonstration.